Distance = rate * time
In the equation D = r * t, to solve for "r", one should:
: Distance = rate * time
In the equation D = r * t, to solve for "r", one should:

Answer by nerdybill(1052)

Just divide both sides by 't'.
.
If you started with:
d = rt
Now, divide both sides by t:
d/t = (rt)/t
Notice the t's on the right cancel:
d/t = r